# CroSTAta: Cross-State Transition Attention Transformer for Robotic Manipulation

### v1

@@ Abstract
Learning robotic manipulation policies through supervised learning from demonstrations remains challenging when policies encounter execution variations not explicitly covered during training. While incorporating historical context through attention mechanisms can improve robustness, standard approaches process all past states in a sequence without explicitly modeling the temporal structure that demonstrations may include, such as failure and recovery patterns. We propose a Cross-State Transition Attention Transformer that employs a novel State Transition Attention (STA) mechanism to modulate standard attention weights based on learned state evolution patterns, enabling policies to better adapt their behavior based on execution history. Our approach combines this structured attention with temporal masking during training, where visual information is randomly removed from recent timesteps to encourage temporal reasoning from historical context. Evaluation in simulation shows that STA consistently outperforms standard cross-attention and temporal modeling approaches like TCN and LSTM networks across all tasks, achieving more than 2× improvement over cross-attention on precision-critical tasks. The source code and data can be accessed at the following link: https://github.com/iit-DLSLab/croSTAta
@@ I Introduction
Imitation learning (IL) has emerged as a promising paradigm for training robotic policies by leveraging expert demonstrations rather than learning a policy from scratch through extensive interaction with the environment [1]. The appeal of IL lies in its data efficiency and ability to leverage human expertise, making it particularly attractive for complex manipulation tasks [2, 3, 4, 5, 6]. However, a fundamental limitation of IL approaches lies in the inherent dependence on the statistical distribution of training data, leading to brittle policies that struggle to handle situations not explicitly observed during training [1, 7, 8]. This becomes even more relevant when deploying these models in unstructured and real-world scenarios where environmental conditions, object properties, or execution dynamics may differ from those observed in demonstrations [7, 9, 10].
@@
However, while enriching demonstrations with failure and recovery examples provides a valuable training signal, simply adding more demonstrations is not a scalable solution. In general, it would be extremely difficult to collect examples for every possible failure scenario. This fundamental limitation highlights the need for approaches that can better leverage the underlying causal dependencies present in the data beyond straightforward sequence imitation. These dependencies span multiple aspects: logical dependencies where low-level actions depend on high-level plans; spatial dependencies where end-effector orientation depends on position; and crucially, temporal dependencies where future actions depend on past execution history. Alternative approaches based on planning [19] and hierarchical [20] policy architectures address some of these challenges with promising results, but the fundamental question of how to extract dependency concepts from data and effectively model them in a policy remains central to achieving robust and adaptive behavior.
@@
The temporal modeling challenge is particularly relevant because many robotic tasks are inherently non-Markovian: appropriate action selection often depends not only on current observations, but also on past observations and actions [5, 21, 22]. For example, manipulation scenarios where the robotic arm occludes critical scene information [23], or multi-stage tasks where early execution steps inform later strategies [3, 24]. In these cases, information from past actions and decisions – such as speed, trajectory curvature, or overall strategy – fundamentally determines the execution of future actions. Despite the importance of historical context, learning long-context robotic policies through imitation learning remains challenging due to spurious correlations in extended observation histories [25].
@@
Current sequence modeling approaches in robotics predominantly treat all temporal elements equally, learning relationships between past and present states primarily through the statistical co-occurrence of elements in demonstrated trajectories [26]. While this approach has shown success in various domains, it may not optimally exploit the structured temporal dependencies in rich demonstrations, where specific past states directly inform corrective actions; thus, more targeted attention mechanisms could better capture these state transition relationships.
@@
We propose a state transition attention mechanism that shifts attention-based temporal processing to focus specifically on how the past informs current action selection. Rather than extracting information from past timesteps and learning how to weight attention across the temporal dimension, our approach directly learns to act based on state transition patterns. This allows policies to leverage historical context by matching current situations to learned temporal patterns during action selection. We evaluate our approach against standard attention mechanism for temporal modeling and demonstrate its particular effectiveness in learning from recovery-rich demonstrations. Moreover, through analysis and ablations of the proposed method, we provide insights into how historical information is retrieved during execution phases, demonstrating that our structured attention mechanisms designed for state transition modeling can significantly enhance policy robustness in sequential decision-making.
The main contributions of the paper are:
@@
  • State Transition Attention (STA), a novel attention mechanism designed to capture state transition patterns in robotic manipulation that, to the best of our knowledge, is the first approach to explicitly modulate attention weights based on learned state evolution patterns for temporal reasoning in manipulation policies;
  • Comprehensive empirical evaluation and ablation studies across four manipulation tasks demonstrating that STA achieves consistent improvements over the standard cross-attention approach (up to 2× on precision-critical tasks) and outperforms traditional temporal modeling methods including TCN and LSTM networks. Our analysis highlights the method’s capability to differentiate between different types of historical context during policy execution.
@@ II Related Work
Figure 2: Graphical representation of how cross-attention works within the Transformer architecture adopted. Here, the query (Q) tokens represent joint values, while the keys (K), values (V) and state (S) tokens encode the overall system information. On the left, a) illustrates standard cross-attention and, on the right, b) depicts State Transition Attention.
@@ II-A Temporal Modeling and Attention Mechanisms
The importance of historical context in robotic tasks has motivated various temporal modeling approaches. Classical methods employ Temporal Convolutional Networks (TCNs) [27] and Long Short-Term Memory (LSTMs) [28] modules for sequence processing, while more sophisticated approaches like dynamic neural advection models incorporate temporal skip connections to handle occlusion through context conditioning on previous observations [29]. The L-MAP framework addresses temporal sequence modeling by learning temporally extended macro-actions for scalable decision-making [30]. Recent transformer-based [31] approaches rely on sequential modeling capabilities for temporal decision-making in robotics. Decision transformers and trajectory transformers apply sequence modeling within reinforcement learning frameworks, modeling task execution as autoregressive sequence prediction [32, 33]. In contrast, imitation learning approaches employ sequence modeling through different strategies: methods like VIMA demonstrate capabilities for long-horizon tasks and generalization [4], while others predict sequences of actions in chunks [3, 22, 34, 35]. In-context learning approaches leverage Transformers’ sequential nature by enriching input sequences with example trajectories that attention mechanisms can retrieve from [36]. Standard attention mechanisms have proven excellent for sequence modeling applications, demonstrating strong noise resistance and long-sequence information retrieval capabilities [37]. Nonetheless, they primarily learn state relationships through statistical co-occurrence rather than explicitly modeling the implications of state evolution. This limitation can lead to failures when execution variations or world state interpretations diverge from training distributions. Recent work has begun addressing this limitation through bidirectional approaches that enforce coherence between past and future predictions [25, 24], yet the fundamental challenge of effectively leveraging structured temporal dependencies in recovery-rich demonstrations remains largely unexplored.
@@ III Method
Effective utilization of historical information becomes critical in robotic manipulation tasks where current observations may be insufficient due to occlusion, ambiguous scene configurations, or partial observability. To achieve this, we need both a mechanism capable of modeling how past states relate to current conditions and training data with informative temporal patterns that can aid learning. Our preliminary tests demonstrate that recovery-rich demonstrations—which contain explicit failure-to-recovery patterns—consistently improve performance across different temporal modeling methods compared to training on successful trajectories alone (Fig. 1). This suggests that such demonstrations provide the necessary informative temporal structures for learning robust policies. Building on this insight, we propose a state transition attention mechanism designed specifically to exploit these temporal patterns.
@@ III-A State Transition Attention Mechanism
We propose a novel modification to standard attention mechanisms that focuses on state transition patterns rather than individual past states. Our approach is motivated by the intuition that the relational patterns, most relevant for current decision-making, emerge from understanding how states evolve over time, particularly, observing direct relationships between subsequent states.
Following an architectural approach similar to [35], we adopt an encoder-decoder structure employing cross-attention to relate action representations in the decoder to state information from the encoder. For example, in the case of robotics, this could be relating planned joint movements (actions) to visual observations and sensor readings (states). Standard cross-attention mechanisms in this setup learn to relate current actions to all present and past states through learned linear projections. These projections must learn representations valid across all temporal distances, while positional embeddings distinguish different timesteps and softmax operations weight the relevance of distant past events. This approach places a significant burden on the attention mechanism to both learn appropriate representations and correctly dampen irrelevant historical information through softmax normalization.
@@
$Softmax(\frac{diag(Q_{t-k:t}K^{T}_{t-k:t})(S_{t-k:t}S^{T}_{t})}{\sqrt{d_{K}d_{S}k}})V_{t}$ (2)
where $Q,K,V,S$ represent independent linear neural network projections of decoder ($Q$) and encoder ($K,V,S$) tokens, subscripts indicate the elements of the sequence from time $t$ back to $k$ elements in the past, and finally, $d_{K}$ and $d_{S}$ refer to the dimension of the relative variable projection. The state transition projection $S$ learns to identify which historical states are most relevant given the current state, creating transition-aware attention values that are then multiplied with the diagonal elements of the $QK^{T}$ matrix, i.e., $Q_{t-k}K^{T}_{t-k},\dots,Q_{t}K^{T}_{t}$.
This operation introduces an additional projection matrix but maintains computational efficiency at inference through standard caching strategies [38] for Q, K and S while restricting the expensive softmax computation to include only recent timesteps rather than the entire sequence. Additionally, since attention scores between $Q$ and $K$ are no longer processed based on temporal position, we incorporate positional embeddings directly into the transition projection output, allowing the model to maintain temporal awareness while focusing on state evolution patterns. A graphical comparison between attention strategies is provided in Fig. 2.
@@ III-B Architecture Design
Building upon the STA component, we design our decoder with standard Transformer blocks, each comprising cross-attention (implemented with STA) and self-attention modules followed by a feed-forward network. This decoder processes input tokens representing robot actions that encompass coordinated motion across all joints. To better exploit relational patterns through self-attention mechanisms, we use multiple input tokens per single robot action, initialized with proprioceptive information and combined with absolute positional embeddings [31] to establish intra-relationships between them. After being processed by the Transformer blocks, the output tokens are individually processed through a multi-layer perceptron (MLP) to produce the target action for each joint, then executed through an underlying PD controller. We refer to this decoder architecture as STA Transformer for brevity in the remainder of the paper.
The encoder deals with world state information processing through convolutional neural networks (CNN) and MLP modules that handle visual and proprioceptive inputs, respectively. The concatenated output of these modules represents the state tokens for the decoder’s cross-attention component. This architecture allows the model to process the evolution of the world state (through cross-attention with STA) and to model the robot’s internal kinematics (through self-attention), providing both stages with tokens of current and past timesteps. A complete overview of the architecture is provided in Fig. 3.
@@ III-C Training Strategy with Temporal Masking
For training, we sample sequences of 16 timesteps where each one uses the previous steps as contextual history, while future information remains masked. We optimize the model using mean squared error (MSE) loss between predicted and ground-truth actions. Additionally, to incentivize information retrieval from historical context and enhance learning of temporal dependencies, we propose a temporal masking strategy applied to visual inputs provided to the encoder. This approach removes all exteroceptive information available at current timesteps, forcing the model to rely on historical context for decision-making. We apply this masking to $k$ consecutive timesteps in each sampled sequence (excluding the first/oldest timestep), where $k$ is randomly sampled from [2, LL/2] with $L$ being the total sequence length. This masking strategy serves a dual purpose: it prevents the model from developing over-reliance on current visual information while simultaneously encouraging the development of robust temporal reasoning capabilities over information available in the history.
@@ IV-A Evaluation Setup
To evaluate the effectiveness of our STA mechanism, we design experiments using simulated manipulation tasks where policies can potentially benefit from historical context during execution. We evaluate our method on four tasks from the ManiSkill benchmark [14] selected for their distinct failure modes: StackCube (Fig. 4a) and PegInsertionSide (Fig. 4b) require precise manipulation; TwoRobotStackCube (Fig. 4c) demands coordinated manipulation of two objects simultaneously where failure in either subtask compromises the overall task; and UnitreeG1TransportBox (Fig. 4d) requires complex coordination involving synchronized torso rotation and arm movements. These tasks feature short horizons (less than 5 seconds) but remain failure-prone due to their demanding execution requirements, making them ideal testbeds for evaluating temporal reasoning capabilities under randomly generated conditions at inference time. Each task uses one or two cameras attached to the workspace table and the robotic manipulator to provide visual information, while proprioceptive state information consists exclusively of joint position values. Action prediction consists in joint position delta values across all evaluated methods. Additional environment details are provided in Fig. 4.
@@ IV-B Data Collection
We collect successful demonstrations containing artificially introduced failure sequences followed by natural recovery behaviors, similar to DAgger methodology [1]. Our data generation process employs policies trained with privileged information using the Proximal Policy Optimization (PPO) algorithm. During trajectory collection, we artificially introduce failure patterns by sampling noise at random execution moments to modify the state representation of world elements for $n$ consecutive steps, forcing the policy to perform suboptimal actions (e.g., misperceiving object positions leading to failed grasps or collisions). Trained to be noise-robust, these policies naturally exhibit recovery behaviours towards the real target after the expiration of the $n$ steps, thus creating demonstrations that contain explicit failure-correction patterns. Crucially, while we artificially induce failures, we label the collected trejectory steps to train exclusively on noise-free action predictions. Although predictions for failure-induced steps do not contribute to the training loss, the visited states and generated trajectory segments remain key components of the sequential network input, providing rich temporal context for learning state transition patterns. We generated 1000 episodes per task for training.

### v2

@@ Abstract
Learning robotic manipulation policies through supervised learning from demonstrations remains challenging when policies encounter execution variations not explicitly covered during training. While incorporating historical context through attention mechanisms can improve robustness, standard approaches process all past states in a sequence without explicitly modeling the temporal structure that demonstrations may include, such as failure and recovery patterns. We propose a Cross-State Transition Attention Transformer that employs a novel State Transition Attention (STA) mechanism to modulate standard attention weights based on learned state evolution patterns, enabling policies to better adapt their behavior based on execution history. Our approach combines this structured attention with temporal masking during training, where visual information is randomly removed from recent timesteps to encourage temporal reasoning from historical context. Evaluation in simulation shows that STA consistently outperforms standard attention approach and temporal modeling methods like TCN and LSTM networks, achieving more than 2× improvement over cross-attention on precision-critical tasks. The source code and data can be accessed at the following link: https://github.com/iit-DLSLab/croSTAta
@@ I Introduction
Imitation learning (IL) has emerged as a promising paradigm for training robotic policies by leveraging expert demonstrations rather than learning a policy from scratch through extensive interaction with the environment [5]. The appeal of IL lies in its data efficiency and ability to leverage human expertise, making it particularly attractive for complex manipulation tasks [29, 3, 15, 18, 37]. However, a fundamental limitation of IL approaches lies in the inherent dependence on the statistical distribution of training data, leading to brittle policies that struggle to handle situations not explicitly observed during training [5, 4, 31]. This becomes even more relevant when deploying these models in unstructured and real-world scenarios where environmental conditions, object properties, or execution dynamics may differ from those observed in demonstrations [4, 35, 25, 30].
@@
However, simply enriching demonstrations is not a scalable solution – it would be extremely difficult to collect examples covering every possible failure scenario. This fundamental limitation highlights the need for approaches that can better leverage the underlying causal dependencies present in the data beyond straightforward sequence imitation. These dependencies span multiple aspects: logical dependencies where low-level actions depend on high-level plans; spatial dependencies where end-effector orientation depends on position; and crucially, temporal dependencies where future actions depend on past execution history. Alternative approaches based on planning [7] and hierarchical [19] policy architectures address some of these challenges with promising results, but the fundamental question of how to extract dependency concepts from data and effectively model them in a policy remains central to achieving robust and adaptive behavior.
@@
The temporal modeling challenge is particularly relevant because many robotic tasks are inherently non-Markovian: action selection often depends not only on the present, but also on past observations and actions [18, 24, 38]. For example, manipulation scenarios where the robotic arm occludes critical scene information [27], or multi-stage tasks where early steps inform later strategies [3, 21]. In these cases, information used for decisions (e.g. speed, trajectory curvature, strategy) can fundamentally determine the execution of future actions. Yet learning long-context robotic policies through imitation learning remains challenging due to spurious correlations in extended observation histories [33].
@@
Current sequence modeling approaches in robotics predominantly treat all temporal elements equally, learning relationships between past and present primarily through statistical co-occurrence of elements in demonstrated trajectories [36]. While this approach has shown success in various domains, it may not optimally exploit the structured temporal dependencies in rich demonstrations, where specific past states inform corrective actions; thus, more targeted attention mechanisms could better capture these state transition relationships.
@@
We propose a state transition attention mechanism that shifts attention-based temporal processing on how the past informs current action selection. Rather than extracting information from past timesteps and learning how to weight attention across the temporal dimension our approach directly learns to act based on state transition patterns. This allows policies to leverage historical context by matching current situations to learned temporal patterns during action selection. We evaluate our approach against standard mechanisms for temporal modeling and demonstrate its particular effectiveness in learning from recovery-rich demonstrations. Moreover, through analysis and ablations of the proposed method, we provide insights into how historical information is retrieved during execution phases, demonstrating that our structured attention mechanisms designed for state transition modeling can significantly enhance policy robustness in sequential decision-making.
The main contributions of this paper are:
@@
  • State Transition Attention (STA), a novel attention mechanism that modulates standard attention weights based on learned state evolution patterns, enabling explicit temporal reasoning over execution history in manipulation policies;
  • Empirical evaluation across four manipulation tasks demonstrating competitive performance of STA over standard attention approaches (up to 2× on precision-critical tasks) and over established temporal modeling baselines including TCN and LSTM, with ablation studies and attention pattern analysis providing insight into the mechanism’s behavior.
@@ II Related Work
Figure 2: Graphical representation of how cross-attention works for computation at time $t$ within the Transformer architecture adopted. Here, the query (Q) tokens represent joint values, while the keys (K), values (V) and state (S) tokens encode the overall system information. On the left, a) illustrates standard cross-attention and, on the right, b) depicts State Transition Attention.
@@ II-A Temporal Modeling and Attention Mechanisms
The importance of historical context in robotic tasks has motivated various temporal modeling approaches. Classical methods employ Temporal Convolutional Networks (TCNs) [17] and Long Short-Term Memory (LSTMs) [11] modules for sequence processing, while more sophisticated approaches like dynamic neural advection models incorporate temporal skip connections to handle occlusion through context conditioning on previous observations [6]. The L-MAP framework addresses temporal sequence modeling by learning temporally extended macro-actions for scalable decision-making [22]. Recent transformer-based [32] approaches rely on sequential modeling capabilities for temporal decision-making in robotics. Decision Transformers and Trajectory Transformers apply sequence modeling within reinforcement learning frameworks, modeling task execution as autoregressive sequence prediction [2, 14]. In contrast, imitation learning approaches employ sequence modeling through different strategies: methods like VIMA demonstrate capabilities for long-horizon tasks and generalization [15], while others predict sequences of actions in chunks [3, 38, 1, 28]. In-context learning approaches [8] leverage Transformers’ sequential nature by enriching input sequences with example trajectories that attention mechanisms can retrieve from. Standard attention mechanisms have proven excellent for sequence modeling applications, demonstrating strong noise resistance and long-sequence information retrieval capabilities [34]. Nonetheless, they primarily learn state relationships through statistical co-occurrence rather than explicitly modeling the implications of state evolution. This limitation can lead to failures when execution variations or world state interpretations diverge from training distributions. Recent work has begun addressing this limitation through bidirectional approaches that enforce coherence between past and future predictions [33, 21], yet the fundamental challenge of effectively leveraging structured temporal dependencies in demonstrations remains largely unexplored.
@@ III Method
Effective utilization of historical information becomes critical in robotic manipulation tasks where current observations alone may be insufficient for reliable action selection, whether due to ambiguous scene configurations, execution imprecisions, or dynamic environments where conditions evolve in ways not fully captured at a single timestep. To achieve this, we need both a mechanism capable of modeling how past states relate to current conditions and training data with informative temporal patterns that can aid learning. Our preliminary tests demonstrate that recovery-rich demonstrations – which contain explicit failure-to-recovery patterns – consistently improve performance across different temporal modeling methods compared to training on successful trajectories alone (Fig. 1). This suggests that such demonstrations provide the necessary informative temporal structures for learning robust policies. Building on this insight, we propose a state transition attention mechanism designed to exploit these temporal patterns.
@@ III-A State Transition Attention Mechanism
We propose a modification to standard attention mechanisms that focuses on state transition patterns rather than individual past states. The key intuition is that the relational patterns most relevant for current decision-making, emerge from understanding how states evolve over time, particularly, observing direct relationships between subsequent states.
Following an architectural approach similar to [28], we use an encoder-decoder with cross-attention to relate decoder actions to encoder state information e.g., in a robotic context, joint movements (actions) to visual/sensor inputs (states). Standard cross-attention mechanisms in this setup learn to relate current actions to all present and past states through learned linear projections. These projections must learn representations valid across all temporal distances, while positional embeddings distinguish different timesteps and softmax operations weight the relevance of historical events. This approach places a significant burden on the attention mechanism to both learn appropriate representations and correctly dampen irrelevant historical information through softmax normalization.
@@
$Softmax(\frac{diag(Q_{t-k:t}K^{T}_{t-k:t})(S_{t-k:t}S^{{\color[rgb]{0,0,0}T}}_{t})}{\sqrt{d_{K}d_{S}k}})V_{t}$ (2)
where $Q,K,V,S$ are independent linear neural network projections of decoder ($Q$) and encoder ($K,V,S$) tokens; subscripts span the sequence from time $t$ back $k$ steps; and $d_{K}$ and $d_{S}$ refer to the dimensions of the respective projection matrices. The state transition projection $S$ learns to identify which historical states are most relevant given the current state, creating transition-aware attention values that are then multiplied with the diagonal elements of $QK^{T}$, i.e., $Q_{t-k}K^{T}_{t-k},\dots,Q_{t}K^{T}_{t}$, deliberately decoupling per-timestep action-state alignment from cross-temporal relevance, which is instead captured by the state transition projection $S$. The normalization factor follows standard scaled dot-product attention practices, accounting for the dimensionality of components [32]. Crucially, the softmax operation in STA is applied only over current timestep tokens ($n$ tokens) rather than across the entire history ($(k+1)\cdot n$ tokens), reducing the computational cost of the exponential operations. However, this is offset by the additional projection $S$ and dot product $S_{t-k:t}S^{T}_{t}$ of $O(kn^{2}d_{S})$, resulting in similar overall computational cost while providing different representational capabilities. At inference, both approaches benefit from caching strategies [12], with STA caching $Q$, $K$, and $S$ projections (with $Q$ and $K$ potentially cached in form of attention scores since they don’t depend on future steps), instead of $K$ and $V$. To maintain temporal awareness while focusing on state evolution patterns, we add element-wise learned absolute positional embeddings into the state transition projection output $S$. A graphical comparison between attention strategies is provided in Fig. 2.
@@ III-B Architecture Design
Building upon the STA component, our decoder uses standard Transformer blocks with STA cross-attention, self-attention, and feed-forward layers. It processes input tokens representing coordinated joint actions. To better exploit relational patterns through self-attention mechanisms, we use an input token per single joint action, initialized with proprioceptive information and absolute positional embeddings [32] to establish intra-relationships between them. Output tokens are individually processed through an MLP to produce the target action for each joint, then executed through an underlying PD controller. We refer to this architecture as STA Transformer for brevity in the remainder of the paper.
The encoder deals with world state information processed through a convolutional neural network (CNN) and an MLP network to handle visual and proprioceptive inputs, respectively. The concatenated output of these modules represents the state tokens for the decoder’s cross-attention layers. The overall architecture allows to process the evolution of the world state (through STA cross-attention) and to model the robot’s internal kinematics (through self-attention), providing both stages with tokens of current and past timesteps. A complete overview of the architecture is provided in Fig. 3.
@@ III-C Training Strategy with Temporal Masking
For training, we sample sequences of 16 timesteps where each one uses the previous steps as contextual history, while future information remains masked. We optimize the model using mean squared error loss between predicted and ground-truth actions. Additionally, to incentivize information retrieval from historical context and enhance learning of temporal dependencies, we propose a temporal masking strategy applied to visual inputs provided to the encoder. All exteroceptive information is removed for $k$ consecutive timesteps (excluding the first/oldest), where $k$ is randomly sampled from [2, LL/2] with LL=16 being the total sequence length. This masking strategy serves a dual purpose: it prevents the model from developing over-reliance on current visual information while simultaneously encouraging the development of robust temporal reasoning capabilities, forcing the model to rely on historical context for decision-making.
@@ IV-A Evaluation Setup
We evaluate STA mechanism on four ManiSkill tasks [26] chosen for distinct dynamics and failure modes (Fig. 4): StackCube and PegInsertionSide require precise manipulation; TwoRobotStackCube demands coordinated manipulation of two objects where failure in either subtask compromises the task; and UnitreeG1TransportBox requires synchronized torso rotation and arm movements. These tasks feature short horizons (less than 5 seconds) but remain failure-prone due to their demanding execution requirements, with task configurations that can cause camera occlusions interfering with visual information, e.g., during the manipulation of pegs in PegInsertionSide and the mutual interference of arms in TwoRobotStackCube, making them ideal testbeds for evaluating temporal reasoning capabilities under randomized inference conditions. Each task uses one or two cameras to capture visual information (see Fig. 4), while proprioceptive state information consists exclusively of joint position values. Action prediction consists of joint position delta values across all evaluated methods.
@@ IV-B Data Collection
We collect demonstrations containing artificially induced failure sequences followed by natural recovery behaviors, following a DAgger-like methodology [5]. Data is generated using policies trained with privileged information via PPO, with noise-robustness explicitly enforced through output noise injection during training. At collection time, random perturbations to the state representation of world elements for $n$ consecutive steps force the policy to perform suboptimal actions (e.g., misperceiving object positions leading to failed grasps or collisions), after which it naturally recover toward the true target, yielding trajectories with explicit failure-correction patterns. Crucially, since failures are artificially induced, we label the collected trajectory steps to train exclusively on noise-free action predictions. Although predictions for failure-induced steps do not contribute to the training loss, the visited states and generated trajectory segments remain key components of the sequential network input, providing rich temporal context for learning state transition patterns. We generated 1000 episodes per task for training.